Hi. I may have a solution but it may not be correct.
I need to give the exact and approximate solutions to 3 decimal places.
(x-12)^2 - 12;
My steps:
x-12= sqrt12; x-12 = - sqrt12
exact solutions are 12+2 = sqrt 3 12-2 = sqrt 3.
Approximate solutions are 15.464, -8.536

Answer by Alan3354(69443)

I need to give the exact and approximate solutions to 3 decimal places.
(x-12)^2 - 12;
My steps:
x-12= sqrt12; x-12 = - sqrt12
exact solutions are 12+2 = sqrt 3 12-2 = sqrt 3.
x = 12 ± sqrt(12)
x = 12 ± 2sqrt(3)
I'm not sure if that's what you have.
-------------------------
x =~ 15.464, 8.536 (not negative 8.536)
--------------------
Approximate solutions are 15.464, -8.536